please help me solve this: a. every whole number greater thsn 4 and less than 20 is the sum of two or more consecutive whole numbers.
b. every whole number between 25 and 50 is the product of two whole numbers greater than 1.

Answer by Edwin McCravy(6938)

please help me solve this: a. every whole number greater thsn 4 and less than 20 is the sum of two or more consecutive whole numbers.

 
a. every whole number greater thsn 4 and less than
   20 is the sum of two or more consecutive whole 
   numbers.

The correct answer is that it is false because 8 cannot 
be written that way.
 
8 can't be the sum of two consecutive integers because
 
1+2=3, 2+3=5, 3+4=7, 4+5=9, every other pair of consecutive
integers will exceed 8
 
8 can't be the sum of three consecutive integers because
 
1+2+3=6, 2+3+4=9, every other group of 3 consecutive ingers 
will exceed 8
 
8 can't be the sum of four consecutive integers because
 
1+2+3+4=10, and every other group of 4 consecutive ingers 
will exceed 8
 
And of course there is no use to try 5 or more consecutive 
integers because even the smallest group of 4 have sum 
greater than 8.
 
Therefore it is false because not every whole number greater 
than 4 and less than 20 is the sum of two or more consecutive 
numbers.  8 is a counter-example.
 

b. every whole number between 25 and 50 is the product of two whole numbers greater than 1.

That is false because 29 is a prime number between 
25 and 50 and 29 can only be written as a product 
of two whole numbers, 29x1.

Edwin